# MADNESS Deblender

### v1

@@ 3.4 VAE-deblender
Figure 7: Schematic of the training of the VAE-deblender using two copies of the encoder network illustrated in Figure 3. The trainable network learns to predict the latent space representation of the central galaxy by minimizing the Euclidian distance (marked in red) between its prediction ($z_{d}$) and the representation of the isolated galaxy ($z_{i}$) obtained by the fixed encoder trained as described in Section 3.2.
@@
Once we have the NF and trained VAE, it is possible to generate new galaxies by sampling from the distribution learned from the normalizing flow and feeding it to the decoder. The goal, hereafter, is to obtain the MAP solution in the latent space with gradient descent. To reduce the time taken for optimization, it is essential to provide the model with reasonable initialization of latent points.
For this, we adopt the idea presented by Arcelin et al. (2020), where the authors retrain only the encoder as a deblender, to find the latent space representation of the central galaxy in postage stamps containing blends. While keeping the pre-trained decoder fixed, the authors provide blended stamps as input to the encoder and compute the reconstruction loss of the decoder output with the isolated central galaxy in the input stamp. Upon minimizing the negative log-likelihood under this setting, the newly retrained encoder outputs the amortized posterior distribution of only the central galaxy in the latent space. In other words, the encoder part of the VAE acts as a deblender, extracting only the central galaxy in the postage stamp. Since the decoder is fixed, this step does not affect the underlying distribution of isolated galaxies and does not require retraining the NF.
Arcelin et al. (2020) observed that the results of their deblender significantly improved upon using non-linear data normalization to deal with the broad range of flux values. In our case, retraining the encoder as a deblender is more difficult because of the choice of linear normalization. So instead of computing the loss in the image space, we shift to the latent space. For this purpose, we start with two sets of the trained encoder as shown in Figure 7. Using the BlendedSims we feed images of an isolated galaxy to the first encoder and to the second encoder we provide a corresponding blended scene and the same central galaxy. Let $z_{i}$ (isolated representation) and $z_{d}$ (deblended representation) be samples from the amortized posterior obtained from the first and second encoders, respectively. Keeping the weights of the first encoder fixed, we retrain only the second encoder to minimize the l2-norm or Euclidian distance between $z_{i}$ and $z_{d}$. We run this training for a maximum of $200$ epochs with the Adam optimizer and a learning rate of $10^{-5}$ along with similar settings for clip value, a learning rate scheduler, and patience as the original VAE.
Once trained, the second network can be combined with the previously trained decoder and used as a deblender similar to Arcelin et al. (2020) as shown in Figure 8. Although this architecture predicts a distribution in the latent space, we concern ourselves with only the mean of this distribution; hereafter, we call this the VAE-deblender. In our algorithm, we use the VAE-deblender only to initialize the latent space before the gradient descent for MAP estimate.
@@ 3.5 Deblending
Figure 9: Schematic of MADNESS workflow. The algorithm takes as input a field of galaxies and their detected positions. The VAE-deblender consisting of convolutional and dense layers initializes the latent space ($Z$), which is followed by the optimization of the sum of the log-prior and the log-likelihood by using the decoder (dense and transposed convolutional layers) to map latent space representations to the image space and the normalizing flow (NF) which uses MAF layers to compute the log-prior. Once the stopping condition is satisfied based on the rate of decrease of the minimization objective, we obtain the maximum a posteriori solution $Z^{*}$ as output. The corresponding image space solutions can be found with a forward pass through the decoder.
The VAE has already proved to be a powerful tool in the domain of generative models. However, it fails to provide an explicit likelihood. By combining the VAE with the flow network, the so-called Flow-VAE can be used to obtain the maximum-a-posteriori solution of the deblending problem. Say we have a field of galaxies $y$, and our goal is to obtain a set of galaxies $X$ such that
@@
where $p(y|X)$ is the data likelihood and $p(X)$ is the prior representing the probability of the reconstructions to look like galaxies. As the prior is hard to evaluate in the pixel space without using symmetry constraints like scarlet, we turn towards the latent space. Since the NF has been trained to model the aggregate posterior of the VAE, we can evaluate the prior term in the latent space using the distribution modeled by the NF. Therefore, Equation 10 is rewritten as
@@
where $Z$ is the set of latent space representations of galaxies, and $\theta$ and $\varphi$ are the pre-trained weights of the VAE and NF, respectively. We can rewrite Equation 11 as the minimization of
@@
The caveat here is that since we have shifted to the latent space for the minimization, we cannot evaluate the log-likelihood directly without mapping to the image space. As the VAE was trained to generate images, the decoder is essentially the mapping back to the image space. To evaluate the log-likelihood term, we use the trained decoder $p_{\theta}$ to go from the latent space representation of each galaxy to the image space as shown in Figure 9. Given the architectural choice of $45\times 45$ pixel cutouts used to train the VAE, the final predicted image of each galaxy is a cutout of the same dimension. By summing over each cutout at the corresponding detection location of the galaxy, we obtain the reconstructed field $\operatorname{\mathbb{X}}$ given by
@@
Assuming uncorrelated Gaussian approximation to Poisson noise in each pixel, we can now compute the log-likelihood term using the reconstructed field as
@@
We use the Adam optimizer for the gradient descent with an initial learning rate of $0.05$ and a scheduler that reduces it by a factor of $0.8$ every $30$ epoch. The optimization is terminated when the decrease in the exponentially weighted moving average of the final $15$ steps drops below a factor of $0.05$ of that of the initial $15$ steps.
Figure 9 summarizes the workflow of the algorithm. After the initialization of the latent space with the VAE-deblender, the decoder allows us to access the first term of Equation 15 while the NF allows us to evaluate the log-prior for each galaxy. For each field, we obtain the MAP solution $\hat{Z}$ in the latent space for all the galaxies simultaneously by minimizing the objective in Equation 15 for all galaxies until the stopping condition is satisfied. Once the solution $\hat{Z}$ is obtained, the corresponding image space solution is obtained by a forward pass through the decoder.
@@ 4 Results
Once we obtain the MAP estimate in the VAE latent space, we use the TestFields described in 2.2 to test the reconstructions in terms of the total flux, g-r color, shapes, and morphology indicators. As a sanity check, we verify that the MAP optimization improves the results from VAE-deblender and we further compare MADNESS against scarlet, using the configuration described in Appendix B.
Although in a more realistic scenario, the deblenders will be used alongside a detection algorithm, here we remove detection biases and report the scores for all the metrics providing the true centers of all galaxies to the deblenders.
@@ 4.1 Metrics
We evaluate the performance of the flux reconstruction using aperture-photometry that computes flux within a given radius around the true center of the galaxies. We compute the g-r color of the galaxies which is the difference in magnitude in the g and r bands, to ensure that the cross-calibration between bands is preserved.
To test the morphology of the reconstructions, we compute SSIM described in Section 3.2 and the cosine similarity which is the cosine of the angle between the reconstructed model and the ground truth flattened as a 1D vector. We characterize the shapes of the galaxies with ellipticities obtained from adaptive moments (see Hirata & Seljak, 2003, for a detailed discussion on adaptive moments), computed using the GalSim HSM module.
The results for these metrics are expected to vary significantly based on the SNR and blending of galaxies. Therefore, to properly understand their impact, we use the following metrics:
@@
    where $S_{n}$ is the vector of true pixel values for the galaxy and $S^{b}_{n}$ is the flux vector from the blended field. For heavily blended objects in the limit of $S^{b}_{n}\gg S_{n}$ where $S_{n}$ is non-zero, $\beta_{n}$ tends to 1. For the other extreme case where the effect of blending is negligible, $S^{b}_{n}\to S_{n}$ near the galaxy so $\beta_{n}$ approaches 0. For a more visual interpretation, Appendix C shows cutouts of galaxies, as the SNR and blendedness vary.
@@
    where $F_{n}^{b}$ and $F_{n}$ are the aperture-photometry measured flux for the blended and isolated galaxy, respectively. Appendix C gives a visual representation of how galaxies look as we vary $\kappa$ in different SNR ranges.
@@
    SNR: In Section 2 a selection cut of $\text{SNR}\geq 10$ was applied on the r-band SNR of each galaxy, computed at the catalog level. For consistency, we continue with the same definition of SNR to evaluate our results.
@@ 4.2 Errors in aperture-photometry of deblended models
(a) Cumulative distribution of the absolute value of relative aperture-photometry flux residual for entire TestFields dataset.
(b) Cumulative distribution of the absolute value of relative aperture-photometry flux residual for blended galaxies with contamination $\kappa_{r}\geq 0.01$.
Figure 10: Comparison of the absolute value of relative residuals $|\delta F/F|$ in aperture-photometry among MADNESS, scarlet, VAE-deblender, and the original field without deblending. The top panel shows the cumulative distribution of the absolute value of the relative residual in the r-band for the entire TestFields dataset while the bottom panel shows the distribution of $|\delta F/F|$ for blended galaxies with $\kappa_{r}>0.01$ which represents at least $1\%$ flux contribution from neighboring sources.
Figure 11: Aperture-photometry results in the $r$-band of galaxies from TestFields after deblending with MADNESS, scarlet, in terms of signal-to-noise ratio (SNR), blendedness ($\beta$), and contamination ($\kappa$) in the r-band. In each panel, the top plot shows the distribution of the absolute value of relative flux residual ($|\delta F/F|$) while the bottom plot shows the distribution of galaxies for the metric on the x-axis. The boxes in the plot on top show the range from the $25^{\rm{th}}$ to $75^{\rm{th}}$ percentile, with a horizontal line representing the median, while the whiskers represent the $5^{\rm{th}}$ to $95^{\rm{th}}$ percentile of relative residuals in the r-band.
For both MADNESS and scarlet, we compute the aperture-photometry for each galaxy by centering on the ground truth position and subtracting the pixel values of all other predicted galaxies from the field except the central one. We choose elliptical apertures with semi-major and semi-minor axis $a$ and $b$ for each band given by
@@
Using the SourceExtractor library (Barbary, 2016) we compute the flux of the reconstructions in each band and plot in Figure 10a the absolute value of relative flux residual which is the ratio of the flux residual ($\delta F$) to the actual flux of the isolated galaxy. Mathematically, the relative flux residual is defined as
@@
where $F^{d}$ and $F$ are the $r$-band aperture-photometry flux in the same noise field for the deblended and isolated galaxy, respectively. To provide a reference for comparison, we also compute the residuals obtained directly on the blended field, i.e., without using a deblender.
For a quantitative comparison, we compute the mean of the absolute value of the relative residuals. However, since the mean can be affected by outliers, we remove severely blended galaxies with contamination $\kappa_{r}>5$ and observe that the mean of the absolute value of the relative residuals in the blended case is $0.142$ which implies on average the residuals are around $14.2\%$ of the flux value. After deblending, we significantly reduce this average to $0.020$ for MADNESS, $0.028$ for scarlet, and $0.039$ for the VAE-deblender. So the average of MADNESS is roughly $29\%$ and $50\%$ lower than that of scarlet and VAE-deblender, respectively.
The cumulative distribution of galaxies as a function of $|\delta F/F|$ in Figure 10a shows that by performing gradient descent in the latent space, MADNESS significantly improves the initial predictions of the VAE-deblender. Moreover, for small values of $|\delta F/F|$, scarlet outperforms MADNESS, but the latter outperforms the former for $|\delta F/F|\gtrapprox 0.005$. The better performance of scarlet for smaller values of $|\delta F/F|$ can be attributed to the fact that scarlet predicts models in the PSF deconvolved state so a large fraction of galaxies that appear as blended for MADNESS are isolated sources for scarlet. To test this hypothesis, we show in Figure 10b, $|\delta F/F|$ for galaxies with $\kappa_{r}\geq 0.01$, i.e., at least $1\%$ flux contribution from neighboring sources, that even for low values of $|\delta F/F|$, MADNESS has a higher density of galaxies than scarlet. This confirms our hypothesis that the better performance of scarlet for small $|\delta F/F|$ is due to very slightly blended galaxies which most likely appear isolated for scarlet after PSF deconvolution.
In Figure 11 we compare the performance of the two deblenders regarding the recovered photometry as a function of SNR, blendedness, and contamination. From these plots, we conclude that MADNESS provides better flux reconstruction than scarlet, especially for heavily blended galaxies; the results are consistent even in different SNR bins.
In Appendix D, we show that our results generalize to other bands except for a deviation in the u-band where MADNESS and VAE-deblender’s performance drops due to the increased noise levels.
@@ 4.3 Bias in color for deblended galaxies
(a) Variation of true color residuals ($\delta C^{T}$) computed with the color from catalog ($C^{T}$) as the baseline.
(b) Distribution of blending error ($\delta C^{B}$) with colors from photometry on noisy isolated galaxies ($C^{N}$) as the baseline.
Figure 12: Comparing the g-r color for galaxies in TestFields for three cases: without deblending, deblending with MADNESS and with scarlet, as we vary the color (top panels) and contamination $\kappa_{r}$ in r-band (bottom panels). For each panel, the top plot shows the color residuals where the boxes and whiskers in the plot indicate the ranges from $25^{\rm{th}}$ to $75^{\rm{th}}$ and $5^{\rm{th}}$ to $95^{\rm{th}}$ percentiles, respectively. The plot at the bottom of each panel shows the distribution of the metric on the horizontal axis.
In Section 4.2 we saw that MADNESS improved the flux reconstruction in the r-band of LSST. However, we must also ensure cross-calibration between bands is conserved as these errors and biases may propagate into downstream analysis, into redshift estimation, for example. In this section, we compare the g-r color obtained with the flux in Section 4.2 for galaxies in TestFields for three cases: without deblending, deblending with MADNESS, and deblending with scarlet. For each of these three cases, we compute the error in measured color against two baselines: the true color from the catalog and the color obtained with photometry on noisy isolated galaxies.

### v2

@@ 3.4 VAE-deblender
Figure 5: Schematic of the training of the VAE-deblender using two copies of the encoder network illustrated in Figure 14. The trainable network learns to predict the latent space representation of the central galaxy by minimizing the Euclidian distance (marked in red) between its prediction ($z_{d}$) and the representation of the isolated galaxy ($z_{i}$) obtained by the fixed encoder trained as described in Appendix B.1.
@@
Once we have the normalizing flow and trained VAE, it is possible to generate new galaxies by sampling from the distribution learned from the normalizing flow and feeding it to the decoder. The goal hereafter is to obtain the MAP solution in the latent space with gradient descent. To reduce the time taken for optimization, it is essential to provide the model with a reasonable initialization of latent points.
For this, we adopt the idea presented by Arcelin et al. (2020), where the authors retrain only the encoder as a deblender, to find the latent space representation of the central galaxy in postage stamps containing blends. While keeping the pre-trained decoder fixed, the authors provide blended stamps as input to the encoder and compute the reconstruction loss of the decoder output with the isolated central galaxy in the input stamp. Upon minimizing the negative log-likelihood under this setting, the newly retrained encoder outputs the amortized posterior distribution of only the central galaxy in the latent space. In other words, the encoder part of the VAE acts as a deblender, extracting only the central galaxy in the postage stamp. Since the decoder is fixed, this step does not affect the underlying distribution of isolated galaxies and does not require retraining the normalizing flow.
Although we follow a similar ideology as Arcelin et al. (2020), the loss function is modified to accommodate our choice of a simpler network architecture and linear data normalization. In our case, retraining the encoder as a deblender is more difficult because of the choice of linear normalization. So, instead of computing the loss in the image space, we shift to the latent space. For this purpose, we start with two sets of the trained encoder as shown in Figure 5. Using the BlendedSims we feed images of an isolated galaxy to the first encoder, and to the second encoder we provide a corresponding blended scene and the same central galaxy. Let $z_{i}$ (isolated representation) and $z_{d}$ (deblended representation) be samples from the amortized posterior obtained from the first and second encoders, respectively. Keeping the weights of the first encoder fixed, we retrain only the second encoder to minimize the L2-norm or Euclidean distance between $z_{i}$ and $z_{d}$ given by
[…]
and the choice of training hyperparameters used is presented in Appendix B.3.
Once trained, the second network can be combined with the previously trained decoder and used as a deblender similar to Arcelin et al. (2020) as shown in Figure 6. Although this architecture predicts a distribution in the latent space, we concern ourselves with only the mean of this distribution; hereafter, we call this the VAE-deblender. In our algorithm, we used the VAE-deblender only to initialize the latent space before the gradient descent for the MAP estimate.
@@ 3.5 Deblending
Figure 7: Schematic of MADNESS workflow. The algorithm takes as input a field of galaxies and their detected positions. The VAE-deblender consisting of convolutional and dense layers initializes the latent space ($Z$), which is followed by the optimization of the sum of the log-prior and the log-likelihood by using the decoder (dense and transposed convolutional layers) to map latent space representations to the image space and the normalizing flow which uses MAF layers to compute the log-prior. Once the stopping condition is satisfied based on the rate of decrease of the minimization objective, we obtain the maximum a posteriori solution $Z^{*}$ as output. The corresponding image space solutions can be found with a forward pass through the decoder.
The VAE has already proved to be a powerful tool in the domain of generative models. However, it fails to provide an explicit likelihood. By combining the VAE with the flow network, the so-called Flow-VAE can be used to obtain the MAP solution of the deblending problem. Say we have a field of galaxies $y$, and our goal is to obtain a set of galaxies $X$ such that
@@
where $p(y|X)$ is the data likelihood and $p(X)$ is the prior representing the probability of the reconstructions to look like galaxies. As the prior is hard to evaluate in the pixel space without using symmetry constraints as in scarlet, we turn towards the latent space. Since the normalizing flow has been trained to model the aggregate posterior of the VAE, we can evaluate the prior term in the latent space using the distribution modeled by the normalizing flow. Therefore, Equation 5 is rewritten as
@@
where $Z$ is the set of latent space representations of galaxies, and $\theta$ and $\varphi$ are the pre-trained weights of the VAE and normalizing flow, respectively. We can rewrite Equation 6 as the minimization of
@@
The caveat here is that since we have shifted to the latent space for the minimization, we cannot evaluate the log-likelihood directly without mapping to the image space. As the VAE was trained to generate images, the decoder is essentially the mapping back to the image space. To evaluate the log-likelihood term, we used the trained decoder $p_{\theta}$ to go from the latent space representation of each galaxy to the image space as shown in Figure 7. Given the architectural choice of $45\times 45$ pixel cutouts used to train the VAE, the final predicted image of each galaxy is a cutout of the same dimension. By summing over each cutout at the corresponding detection location of the galaxy, we obtain the reconstructed field $\operatorname{\mathbb{X}}$ given by
@@
Assuming an uncorrelated Gaussian approximation to Poisson noise in each pixel, we can now compute the log-likelihood term using the reconstructed field as
@@
We used the Adam optimizer for the gradient descent with an initial learning rate of $0.05$ and a scheduler that reduces it by a factor of $0.8$ every $30$ epochs. The optimization is terminated when the decrease in the exponentially weighted moving average of the final $15$ steps drops below a factor of $0.05$ of that of the initial $15$ steps.
Figure 7 summarizes the workflow of the algorithm. After the initialization of the latent space with the VAE-deblender, the decoder allows one to evaluate the first term of Equation 10 while the normalizing flow allows the evaluation of the log-prior for each galaxy. For each field, we obtain the MAP solution $\hat{Z}$ in the latent space for all the galaxies simultaneously by minimizing the objective in Equation 10 for all galaxies until the stopping condition is satisfied. Once the solution $\hat{Z}$ is obtained, the corresponding image space solution is obtained by a forward pass through the decoder.
@@ 4 Results
Once we obtain the MAP estimate in the VAE latent space, we use the TestFields described in 2.2 to test the reconstructions in terms of the total flux, g-r color, shapes, and morphology indicators. As a sanity check, we verify that the MAP optimization improves the results from the VAE-deblender and we further compare MADNESS against scarlet, using the configuration described in Appendix D.
These tests were run on single V100 GPUs at CC-IN2P3, where we parallelized the computation of 20 fields from the TestFields dataset at a time and observed a gradient descent convergence within 1 second per field (approximately 16 galaxies/sec).
Although in a more realistic scenario, the deblenders will be used alongside a detection algorithm, here we remove detection biases and report the scores for all the metrics, providing the true centers of all galaxies to the deblenders.
@@ 4.1 Metrics
We evaluated the performance of the flux reconstruction using aperture-photometry that computes flux within a given radius around the true center of the galaxies. We compute the g-r color of the galaxies, which is the difference in magnitude in the g and r bands, to ensure that the cross-calibration between bands is preserved.
To test the morphology of the reconstructions, we computed the SSIM described in Section 3.2 and the cosine similarity, which is the cosine of the angle between the reconstructed model and the ground truth flattened as a 1D vector. We characterize the shapes of the galaxies with ellipticities obtained from adaptive moments (see Hirata & Seljak, 2003, for a detailed discussion on adaptive moments), computed using the GalSim HSM module.
The results for these metrics are expected to vary significantly based on the S/N and blending of galaxies. Therefore, to properly understand their impact, we used the following metrics:
@@
    where $S_{n}$ is the vector of true pixel values for the galaxy and $S^{b}_{n}$ is the flux vector from the blended field. For heavily blended objects in the limit of $S^{b}_{n}\gg S_{n}$ where $S_{n}$ is non-zero, $\beta_{n}$ tends to 1. For the other extreme case where the effect of blending is negligible, $S^{b}_{n}\to S_{n}$ near the galaxy so $\beta_{n}$ approaches 0. For a more visual interpretation, Appendix E shows cutouts of galaxies, as the S/N and blendedness vary.
@@
    where $F_{n}^{b}$ and $F_{n}$ are the aperture-photometry measured flux for the blended and isolated galaxy, respectively. Appendix E gives a visual representation of how galaxies look as we vary $\kappa$ in different S/N ranges.
@@
    S/N: In Section 2, a selection cut of $\text{S/N}\geq 10$ was applied on the r-band S/N of each galaxy, computed at the catalog level. For consistency, we continue with the same definition of S/N to evaluate our results.
@@ 4.2 Errors in aperture-photometry of deblended models
Figure 8: Comparison of the absolute value of relative residuals $|\delta F/F|$ in aperture-photometry among MADNESS, scarlet, VAE-deblender, and the original field without deblending. Panel (a) shows the cumulative distribution of the absolute value of the relative residual in the r-band for the entire TestFields dataset, while panel (b) shows the distribution of $|\delta F/F|$ for blended galaxies with $\kappa_{r}>0.01$, which represents at least $1\%$ flux contribution from neighboring sources.
Figure 9: Aperture-photometry results in the $r$-band of galaxies from TestFields after deblending with MADNESS, scarlet, in terms of signal-to-noise ratio (S/N), blendedness ($\beta$), and contamination ($\kappa$) in the r-band. In each panel, the top plot shows the distribution of the absolute value of relative flux residual ($|\delta F/F|$) while the bottom plot shows the distribution of galaxies for the metric on the x-axis. The boxes in the plot on top show the range from the $25^{\rm{th}}$ to $75^{\rm{th}}$ percentile, with a horizontal line representing the median, while the whiskers represent the $5^{\rm{th}}$ to $95^{\rm{th}}$ percentile of relative residuals in the r-band.
For both MADNESS and scarlet, we compute the aperture-photometry for each galaxy by centering on the ground truth position and subtracting the pixel values of all other predicted galaxies from the field except the central one. We choose elliptical apertures with semi-major and semi-minor axes $a$ and $b$ for each band given by
@@
Using the SourceExtractor library (Barbary, 2016), we computed the flux of the reconstructions in each band. Figure 8a shows the absolute value of the relative flux residual, which is the ratio of the flux residual ($\delta F$) to the actual flux of the isolated galaxy. Mathematically, the relative flux residual is defined as
@@
where $F^{d}$ and $F$ are the $r$-band aperture-photometry flux in the same noise field for the deblended and isolated galaxy, respectively. To provide a reference for comparison, we also computed the residuals obtained directly on the blended field, i.e., without using a deblender.
For a quantitative comparison, we computed the mean of the absolute value of the relative residuals. However, since the mean can be affected by outliers, we removed severely blended galaxies with contamination $\kappa_{r}>5$ and observed that the mean of the absolute value of the relative residuals in the blended case is $0.142$ which implies on average, the residuals are around $14.2\%$ of the flux value. After deblending, we significantly reduce this average to $0.020$ for MADNESS, $0.028$ for scarlet, and $0.039$ for the VAE-deblender. So the average of MADNESS is roughly $29\%$ and $50\%$ lower than that of scarlet and VAE-deblender, respectively.
The cumulative distribution of galaxies as a function of $|\delta F/F|$ in Figure 8a shows that by performing gradient descent in the latent space, MADNESS significantly improves the initial predictions of the VAE-deblender. Moreover, for small values of $|\delta F/F|$, scarlet outperforms MADNESS, but the latter outperforms the former for $|\delta F/F|\gtrapprox 0.005$. The better performance of scarlet for smaller values of $|\delta F/F|$ can be attributed to the fact that scarlet predicts models in the PSF deconvolved state, so a large fraction of galaxies that appear as blended for MADNESS are isolated sources for scarlet. To test this hypothesis, we show in Figure 8b, $|\delta F/F|$ for galaxies with $\kappa_{r}\geq 0.01$, i.e., at least $1\%$ flux contribution from neighboring sources, that even for low values of $|\delta F/F|$, MADNESS has a higher density of galaxies than scarlet. This confirms our hypothesis that the better performance of scarlet for small $|\delta F/F|$ is due to very slightly blended galaxies, which most likely appear isolated for scarlet after PSF deconvolution.
In Figure 9, we compare the performance of the two deblenders regarding the recovered photometry as a function of S/N, blendedness, and contamination. From these plots, we conclude that MADNESS provides better flux reconstruction than scarlet, especially for heavily blended galaxies; the results are consistent even in different S/N bins.
In Appendix F, we show that our results generalize to other bands except for a deviation in the u-band where MADNESS and VAE-deblender’s performance drops due to the increased noise levels.
@@ 4.3 Bias in color for deblended galaxies
Figure 10: Comparison of the g-r color for galaxies in TestFields for three cases: without deblending, deblending with MADNESS and with scarlet, as we vary the color (top panels) and contamination, $\kappa_{r}$, in the r-band (bottom panels). Panel (a) shows the variation of true color residuals ($\delta C^{T}$) computed with the color from catalog ($C^{T}$) as the baseline, and panel (b) shows the distribution of blending error ($\delta C^{B}$) with colors from photometry on noisy isolated galaxies ($C^{N}$) as the baseline. For each panel, the top plot shows the color residuals where the boxes and whiskers in the plot indicate the ranges from $25^{\rm{th}}$ to $75^{\rm{th}}$ and $5^{\rm{th}}$ to $95^{\rm{th}}$ percentiles, respectively. The plot at the bottom of each panel shows the distribution of the metric on the horizontal axis.
In Section 4.2, we saw that MADNESS improved the flux reconstruction in the r-band of LSST. However, we must also ensure the cross-calibration between bands is conserved, as these errors and biases may propagate into downstream analysis, into redshift estimation, for example. In this section, we compare the g-r color obtained with the flux in Section 4.2 for galaxies in TestFields for three cases: without deblending, deblending with MADNESS, and deblending with scarlet. For each of these three cases, we compute the error in measured color against two baselines: the true color from the catalog and the color obtained with photometry on noisy isolated galaxies.